OpenStudy (anonymous):

How do you convert this to rectangular form?? x = (80cos(35))t y = 7 + (80sin(35))t - 16t^2

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

isolate t in the first equation, then plug it into the second equation

jimthompson5910 (jim_thompson5910):

x = (80cos(35))t x/(80cos(35)) = t t = x/(80cos(35))

jimthompson5910 (jim_thompson5910):

plug that into y = 7 + (80sin(35))t - 16t^2

OpenStudy (anonymous):

Oh, sorry afk

jimthompson5910 (jim_thompson5910):

thats ok

OpenStudy (anonymous):

So it is -((16)x^2)+x+7???

OpenStudy (anonymous):

@ganeshie8

jimthompson5910 (jim_thompson5910):

not quite

OpenStudy (anonymous):

What did I mess up?

jimthompson5910 (jim_thompson5910):

you should have a sin(35)/cos(35) term in there somewhere

jimthompson5910 (jim_thompson5910):

or a tan(35) since sin(x)/cos(x) = tan(x)

OpenStudy (anonymous):

So it is -((16)x^2/tan35)+x+7

jimthompson5910 (jim_thompson5910):

closer but please post your steps so I can see how you got that

OpenStudy (anonymous):

Nevermind, I did it over, and I see.

jimthompson5910 (jim_thompson5910):

alright great